2015-03-10 physics coordinator report » History » Revision 17

Revision 16 (Rogers, Chris, 10 March 2015 14:33) → Revision 17/20 (Rogers, Chris, 10 March 2015 14:35)

h1. physics coordinator report 

 * Step I papers:- 
 ** EMR data paper- VC talk, analysis completed and now in "write up" phase 
 ** PID paper:- hung up on reconstruction issues; fixed unpacking has been run 
    but ADC response of the KL now looks different. Discussion Monday pm 
 ** (Polarisation analysis - held up by dodgy EMR code) 
 * Cooling demo paper: 
 ** Secondary absorber did not present sufficient material to the beamline to 
    shield the tracker, review of secondary absorber position which has caused 
    delay to the paper. 
 ** Tried for a revision this week for review, it didn't work out, hope to have a 
    revision by next week 
 * Step IV papers: 
 ** Status at 
 ** Trying to bring material physics people "out of the woodwork" 
 * Rogers is re-establishing selection/weighting algorithms (dead since 2007) 
 ** aim to drive an analysis of "physics block challenge" before MPB 
 ** feed into review of data rate issues, etc. 
 * Field mapping 
 ** Blackmore/Cobb have a good analysis of the field axis in the FC, which does not  
    produce the "tails" ruining the fitting, hopeful that SS will follow in time 
    for alignment beginning of April  
 * March/April running 
 ** Rogers will be physics shifter on March 21st 
 ** Trying to establish physics shifters on March 28th and following 
 * June running 
 ** Planning note in preparation, due Friday 
 ** Discuss/finalise note at a meeting Thursday 19th 
 ** Tracker will not be able to get plan in time for the meeting 
 ** EMR request one week data and additionally some cosmics 
 ** C+M request one week 
 ** TOF request calibration (1 or 2 shifts) 
 ** KL request cosmic data taking 
 ** Ckov request 1 shift 
 ** Beamline (hardware) request an activation run 
 ** Beamline (optics) not sure 
 ** Cooling channel first data may at some point come through, not sure